A wire has a diameter of 5mm, length 50m, resistivity 3.2 x 10^-8 ohmsm. What is the resistance of the wire?


1
Expert's answer
2021-02-24T16:59:07-0500

We can find the resistance of the wire from the formula:


"R=\rho\dfrac{l}{A},""R=3.2\cdot10^{-8}\ \Omega\cdot m\cdot\dfrac{50\ m}{\pi\cdot\dfrac{(5\cdot10^{-3}\ m)^2}{4}}=0.081\ \Omega."